# Tracing attosecond currents and controlling their direction in a scanning tunneling microscope

### v1

@@ Discussion
Our work shows that two-color sub-cycle waveform control can induce a train of sub-femtosecond current bursts across a variable STM junction and robustly control their direction, allowing for a lateral spatial resolution of 2 nm and the ability to sense sub-angström topographic features under ambient conditions and free from thermal artifacts. In the future, our approach can be combined with all-optical readout 39 to directly reveal the evolution of the sub-cycle currents. Moving from a two-color laser pulse to a single-cycle infrared pulse will naturally lead to a single isolated attosecond electron burst with waveform-controlled timing and direction. Combining the ability to generate these bursts with the well-established capability to image electronic states in molecules, defects, and nanostructures using conventional STM in ultrahigh vacuum conditions, we envision the combined real-time and real-space observations of coherent electron-hole dynamics and many-body effects as they unfold. For example, a sub-cycle controlled infrared pump pulse will enable extracting an electron from a molecule with attosecond precision in time at a specific atomic site, inducing a coherent evolution of complex electron-hole dynamics. A second, sub-cycle controlled probe pulse with opposite field waveform can then inject an electron back into the molecule, allowing for a readout of the dynamics before dephasing effects set in. With attosecond STM, microscopy has the potential to reach its ultimate resolution limits at the atomic scale.
@@ Methods
100 fs laser pulses from an Er:fiber laser system (Menlo C-Fiber High Power, 500 mW, 1550 nm, 80 MHz) are focused into a fiber assembly (standard polarization-maintaining fiber of 75 mm length spliced to 12 mm of highly nonlinear fiber, Thorlabs HN1550P) for supercontinuum generation spanning from 950 nm to 2100 nm (cf. 49). We use only the solitonic part of the resulting supercontinuum with a central wavelength of 1850 nm by rejecting the spectral components below 1500 nm using a dichroic filter. These pulses are focused into a bismuth triborate (BiBO) crystal with a thickness of 1 mm for second harmonic (SH) generation, avoiding the messy part of the spectrum below 1650 nm by an appropriate choice of phase-matching conditions. The recollimated fundamental and SH beams are then fed into an interferometric setup where they are separated by a dispersion-engineered dichroic mirror. The SH beam passes through several filters to remove any remaining trace of the fundamental beam. Subsequently, its polarization is rotated to match the polarization of the fundamental. It then hits a mirror glued onto a piezo chip at a near-zero angle of incidence. The piezo can be modulated at kHz frequencies using a high-voltage signal to periodically change the temporal delay of the SH beam, providing the modulation to produce the lock-in signal. The delay modulation amplitude in our experiment is fs. The fundamental beam is passed through 1 mm of silicon to compensate for the dispersion of the other optical elements in the system and a longpass filter at 1650 nm to remove the messy part of the spectrum below that wavelength. We also control the base two-color delay using a mirror mounted on a linear closed-loop piezo stage at a near-zero angle of incidence. Both beams are recombined by another dispersion-engineered dichroic mirror and expanded by a factor of 3. We monitor the two-color delay using an auxiliary collinear interferometer setup and correct for drifts.
@@
Before sending the beams to the microscope (RHK Technologies PanScan Flow Kit operated in ambient conditions), we control the laser power of the combined beams using a variable neutral-density filter and tune their joint polarization by an achromatic half-wave plate. An off-axis parabolic mirror with a focal length of 50 mm focuses the beams on the tip-sample junction at an angle of incidence of 20 degrees with respect to the sample plane. We use commercial Pt:Ir nanotips (Unisoku P-100PtIr(S)) with apex curvature radii of 5…20 nm and a flat gold substrate fabricated according to the procedure described in 36. The fundamental laser beam is focused to a spot size of 17 m. The duration of the pulse is 35 fs. The SH beam is focused to a spot size of 10 m. The duration of the SH pulse is 62 fs due to uncompensated chirp. Taking into account all parameters, the ratio of the SH intensity to the fundamental intensity is about 10 %. The gold sample can be biased, and the current is measured using a current preamplifier connected to the tip. In the tip freezing experiments, we switch off the bias and the microscope feedback loop for a duration of 600 ms and take several data points during this time.
@@
Our theory modeling is based on the one-dimensional time-dependent Schrödinger equation (TDSE) in the single active electron approximation. Despite the strong degree of simplification, such an approach has yielded excellent agreement with nanotip experiments 19, 4, for instance. Tip and sample are modeled as potential wells and populated with a wavefunction at the Fermi level (see Supplementary Information for more details). We follow two approaches, a numerical integration of the TDSE (nTDSE) based on the Crank-Nicolson method (see Supplementary Information) and an analytical strong-field (SF) model. In the following, we briefly describe the SF model 42, 44.
Based on the Dyson equation, we define the tunneling amplitude from tip to sample as
@@
In analogy to the strong-field approximation of attosecond science, the wavefunctions and can be obtained by using the Van Vleck propagator and the eigenfunction of the sample system 44. The simplified tunneling amplitude is now
Here is a prefactor originating from the transition matrix elements. For an initial bound state in the tip and a final state in the sample, the action in the exponent is given by
@@ Acknowledgements
The authors thank M. Ivanov, A. Kollin and A. Borisov for insightful discussions, I. Kaminer for providing lab space for initial experiments, A. Feigenboim for supporting the design of experimental system, I. Sthzeglowski for the initial setup of the fiber laser, P. Sidorenko for assistance with fiber splicing, and N. Dudovich, U. Leonhardt and L. M. Procopio for providing specialized equipment. This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 853393-ERC-ATTIDA and from the Israel Science Foundation (ISF) under grant agreement No 1504/20. We also acknowledge the Helen Diller Quantum Center and the Russell Berrie Nanotechnology Institute at the Technion for partial financial support.

### v2

@@ Conclusion
Our work shows that two-colour sub-cycle waveform control can induce a train of sub-femtosecond current bursts across a variable STM junction and robustly control their direction, allowing for a lateral spatial resolution of 2 nm and the ability to sense sub-angström topographic features under ambient conditions and free from thermal artifacts. In the future, our approach can be combined with all-optical readout 43 to directly reveal the evolution of the sub-cycle currents. Moving from a two-colour laser pulse to a single-cycle infrared pulse will naturally lead to a single isolated attosecond electron burst with waveform-controlled timing and direction. Combining the ability to generate these bursts with the well-established capability to image electronic states in molecules, defects, and nanostructures using conventional STM in ultrahigh vacuum conditions, we envision the combined real-time and real-space observations of coherent electron-hole dynamics and many-body effects as they unfold. For example, a sub-cycle controlled infrared pump pulse will enable extracting an electron from a molecule with attosecond precision in time at a specific atomic site, inducing a coherent evolution of complex electron-hole dynamics. A second, sub-cycle controlled probe pulse with opposite field waveform can then inject an electron back into the molecule, allowing for a readout of the dynamics before dephasing effects set in. With attosecond STM, microscopy has reached its ultimate spatiotemporal resolution limits at the atomic scale.
@@ Methods
100 fs laser pulses from an Er:fiber laser system (Menlo C-Fiber High Power, 500 mW, 1550 nm, 80 MHz) are focused into a fiber assembly (standard polarization-maintaining fiber of 75 mm length spliced to 12 mm of highly nonlinear fiber, Thorlabs HN1550P) for supercontinuum generation spanning from 950 nm to 2100 nm (cf. 52). We use only the solitonic part of the resulting supercontinuum with a central wavelength of 1850 nm by rejecting the spectral components below 1500 nm using a dichroic filter. These pulses are focused into a bismuth triborate (BiBO) crystal with a thickness of 1 mm for second harmonic (SH) generation, avoiding the messy part of the spectrum below 1650 nm by an appropriate choice of phase-matching conditions. The recollimated fundamental and SH beams are then fed into an interferometric setup where they are separated by a dispersion-engineered dichroic mirror. The SH beam passes through several filters to remove any remaining trace of the fundamental beam. Subsequently, its polarization is rotated to match the polarization of the fundamental. It then hits a mirror glued onto a piezo chip at a near-zero angle of incidence. The piezo can be modulated at kHz frequencies using a high-voltage signal to periodically change the temporal delay of the SH beam, providing the modulation to produce the lock-in signal. The delay modulation amplitude in our experiment is fs. The fundamental beam is passed through 1 mm of silicon to compensate for the dispersion of the other optical elements in the system and a longpass filter at 1650 nm to remove the messy part of the spectrum below that wavelength. We also control the base two-colour delay using a mirror mounted on a linear closed-loop piezo stage at a near-zero angle of incidence. Both beams are recombined by another dispersion-engineered dichroic mirror and expanded by a factor of 3. We monitor the two-colour delay using an auxiliary collinear interferometer setup and correct for drifts.
@@
Before sending the beams to the microscope (RHK Technologies PanScan Flow Kit operated in ambient conditions), we control the laser power of the combined beams using a variable neutral-density filter and tune their joint polarization by an achromatic half-wave plate. An off-axis parabolic mirror with a focal length of 50 mm focuses the beams on the tip-sample junction at an angle of incidence of 20 degrees with respect to the sample plane. We use commercial Pt:Ir nanotips (Unisoku P-100PtIr(S)) with apex curvature radii of 5…20 nm and a flat gold substrate fabricated according to a self-formation procedure 53. The fundamental laser beam is focused to a spot size of 17 m. The duration of the pulse is 35 fs. The SH beam is focused to a spot size of 10 m. The duration of the SH pulse is 62 fs due to uncompensated chirp. Taking into account all parameters, the ratio of the SH intensity to the fundamental intensity is about 10 %. The gold sample can be biased, and the current is measured using a current preamplifier connected to the tip. In the tip freezing experiments, we switch off the bias and the microscope feedback loop for a duration of 600 ms and take several data points during this time.
@@
Our theory modeling is based on the one-dimensional time-dependent Schrödinger equation (TDSE) in the single active electron approximation. Despite the strong degree of simplification, such an approach has yielded excellent agreement with nanotip experiments 13, 19, for instance. Tip and sample are modeled as potential wells and populated with a wavefunction at the Fermi level (see Supplementary Information for more details). We follow two approaches, a numerical integration of the TDSE (nTDSE) based on the Crank-Nicolson method (see Supplementary Information) and an analytical strong-field (SF) model. In the following, we briefly describe the SF model 44, 46.
Based on the Dyson equation, we define the tunnelling amplitude from tip to sample as
@@
In analogy to the strong-field approximation of attosecond science, the wavefunctions and can be obtained by using the Van Vleck propagator and the eigenfunction of the sample system 46. The simplified tunnelling amplitude is now
Here, is a prefactor originating from the transition matrix elements. For an initial bound state in the tip and a final state in the sample, the action in the exponent is given by
@@ Acknowledgements
The authors thank A. Borisov, M. Ivanov, A. Kollin and Y. Bekenstein for insightful discussions, I. Kaminer for providing lab space for initial experiments, A. Feigenboim for supporting the design of experimental system, I. Sthzeglowski for the initial setup of the fiber laser, P. Sidorenko for assistance with fiber splicing, and N. Dudovich, U. Leonhardt and L. M. Procopio for providing specialized equipment. This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 853393-ERC-ATTIDA and from the Israel Science Foundation (ISF) under grant agreement No 1504/20. We also acknowledge the Helen Diller Quantum Center and the Russell Berrie Nanotechnology Institute at the Technion for partial financial support.